OpenStudy (anonymous):

sqrt 21+x + sqrt 21-6 / sqrt 21+6 - sqrt 21-6 = 21/x

OpenStudy (kkutie7):

\[\frac{\sqrt{21+x}+\sqrt{21-6}}{\sqrt{21+6}-\sqrt{21-6}}=\frac{21}{x}\]

OpenStudy (anonymous):

that is the problem what is x?? pls